OpenStudy (anonymous):

Solve the equation 5x = 25

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

You are trying to get x by itself. So all you need to do is divide by the 5. Thus x=5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome(:

OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. (2 + x) + y = 2 + (x + y)

OpenStudy (anonymous):

Commutative Property of Addition

OpenStudy (anonymous):

thanks can you help me with more please

OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. x + 5 = x + 5

OpenStudy (anonymous):

maybe Identity property? not such though sorry

OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. x • y • z = y • x • z Distributive Property Transitive Property Associative Property of Multiplication Commutative Property of Multiplication

OpenStudy (anonymous):

Commutative Property of Multiplication because the variables can be in any order and get the same result.